OpenStudy (anonymous):

-x/3-10=-20 solve for x

OpenStudy (anonymous):

-x-30=-60 x=30

OpenStudy (anonymous):

-x/3-10=-20 -x/3=-10 x/3=10 x=30

OpenStudy (anonymous):

-x/3-10=-20 x/3=10 x=30

OpenStudy (anonymous):

\[\frac{-x}{3-10}=-20?\]

OpenStudy (anonymous):

or: \[\frac{-x}{3}-10=-20?\]

OpenStudy (anonymous):

the second one Nancy Lam

OpenStudy (anonymous):

if the second one than \[\frac{-x}{3}=-20+10\]\[-x=-10*3\]\[x=30\]

OpenStudy (anonymous):

alrighty thanks i got the answer now